OpenStudy (clara1223):

simplify the following, where k represents any integer: cos(2kpi-x)(-sin(2kpi-x)) a) -cosx*sinx b) (1/2)sin(2x) c) cos((1/2)x)^2 d) cos(x)^2*sin(x)^2 e) undefined

OpenStudy (clara1223):

@Loser66 could you explain how you got that answer please?

OpenStudy (mathstudent55):

Here is the explanation: Since the sine and cosine are periodic functions with a period of 2pi, sin (2kpi - x) = sin (-x) and cos(2kpi - x) = cos (-x) Ok so far?

OpenStudy (mathstudent55):

Then because of even and odd functions, sin (-x) = - sin x and cos (-x) = cos x

OpenStudy (mathstudent55):

Putting it all together, you get: \((\cos (2k \pi-x)) ((-\sin(2k \pi-x)) \) \(= \cos (-x) \sin (-x)\) \(= - \sin x \cos x\)

OpenStudy (clara1223):

@mathstudent55 thanks so much!

OpenStudy (dumbcow):

wait ...... -sin(-x) = sin(x) it equals positive sin(x) cos(x) then use double angle identity for sin ----> sin(2x) = 2sincos ----> sincos = 1/2 sin(2x)

OpenStudy (mathstudent55):

@dumbcow You're correct. I missed the negative with the sin in the original problem.

OpenStudy (mathstudent55):

This is how it should read: \((\cos (2k \pi-x)) ((-\sin(2k \pi-x))\) \(= \cos (-x)(- \sin (-x))\) \(= \cos x (-(-\sin x))\)\) \(= \sin x \cos x\) \(=\dfrac{1}{2} \sin 2x\)